OpenStudy (anonymous):

Y=0.04x ^2 + 8.3x y= 1.12x + 1.6 a)no solution b) 0.22, 1.85 c) 179.28, 202.39 d) 0.22, 1.85 and 129.28, 202.39

OpenStudy (anonymous):

I am guessing that you're trying to find the intersection of these two equations. Set the two equal to each other, so .04x^2 + 8.3 x = 1.12x + 1.6 .04x^2 + 7.18x - 1.6 = 0 Find the roots of the quadratic equation above and you get x = -179.2 and x = .2225, those are the two points where the two equations intersect!

OpenStudy (anonymous):

i donno how to do this stuff at all .... on top of its been like 6 years pluys since done this

OpenStudy (anonymous):

Do you understand how I got the quadratic equation: .04x^2 + 7.18x - 1.6 = 0 by setting the original two equations equal to each other?